the measure of two complementary angles are represented by (3x+15) and (2x-10) what is the value of x?
Answer by stanbon(75887)

the measure of two complementary angles are represented by (3x+15) and (2x-10) what is the value of x?
---
If they are complementary, their sum is 90.
---
Equation:
3x+15 + 2x-10 = 90
5x + 5 = 90
5x = 85
x = 17
---
Angle: 3x+15 = 3*17+15 = 66 degrees
Angle 2x-10 = 2*17-10 = 24 degrees
------
Cheers,
Stan H.